OpenStudy (anonymous):

Trinomial 1. 2x^2-5x+2 2. 3x^2+10x+3 3. 2x^2+11x+5 4. 2x^2+7x+6 5. x^2-3x-40 please help me solve i cant figure it out

OpenStudy (anonymous):

How are u supposed to solve them??? By factorising????

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

2x^2-5x+2 STEP 1 : we find the product of the coefficient of x^2 (i.e. 2) and constant (i.e. 2) so 2 * 2 = 4 STEP 2: Now we find two factors of 4, say A and B, such that A*B = 4 and A+B = -5 (i.e. coefficient of x) 4 = 1 x 4 or -1 x -4 4 = 2 x 2 or -2 x -2 we find -1 * -4 = 4 and (-1) + (-4) = -5 STEP 3: Now we split the middle term -5x as -x -4x so equation becomes 2x^2 -x -4x + 2 x(2x - 1) -2(2x - 1) (x - 2)(2x -1)

OpenStudy (anonymous):

like for example 2x^2-7x-15 (2x+3)(x-5) answer

OpenStudy (anonymous):

Now apply the same method on all others and you will get your answers....

OpenStudy (anonymous):

Yes it is correct...